# Information-Theoretic Thresholds for Bipartite Latent-Space Graphs Under Noisy Observations

### v1

@@ 2.4. Bounding expected signed weights
While an extension of the bounds given in the recent work of Bangachev and Bresler [3] to large $\alpha$ would be no easy task on its own, it turns out that, even if this was possible and the bounds given in [3] would extend as they are to our setting, they would nonetheless be far too weak for our purposes, even in the noiseless case. The reason for this is that (as described in [3, Remark 1]) the bounds given in [3] are of order at least $(1/\sqrt{d})^{|V(\alpha)|}$, which is not enough to ensure a small sum in (2.1). The reason for this is that a decay of order $(1/\sqrt{d})^{|V(\alpha)|}$ fails to counterbalance the number of possible $\alpha\subseteq K_{n,m}$ already if $\alpha$ has only $n^{\varepsilon}$ many vertices. To improve upon this, we would instead require bounds that decay exponentially in $|\alpha|$ (i.e. the number of edges) instead of the number of vertices $|V(\alpha)|$.
One possible approach for dealing with this issue (as also suggested in [3, Remark 1]) is to condition on a subset of the latent information. While this is indeed an important aspect of our approach, the problem with the arguments given in [3] is that, even after conditioning and switching to our bipartite setting, we would still not get an exponent that is strong enough for our purposes. Moreover, as we will see later, we do not only need a bound on $\mathbb{E}\left[\textsc{Sw}(\alpha)\right]$, but (as we will see later), we also need to make an explicit statement about how exactly this quantity depends on the latent vectors we condition on. This does not seem to be within reach of techniques considered so far.

### v3

@@ 2.4. Bounding expected signed weights
While an extension of the bounds given in the recent work of Bangachev and Bresler [3] to large $\alpha$ would be no easy task on its own, it turns out that even if the bounds given in [3] would extend to larger $\alpha$ as they are, they would nonetheless be too weak for our purposes, even in the noiseless case. The reason for this is that (as described in [3, Remark 1]) the bounds given in [3] are of order at least $(1/\sqrt{d})^{|V(\alpha)|}$, while we would need a decay of order $(1/\sqrt{d})^{|\alpha|}$ (i.e. a bound that decays exponentially in the number of edges instead of vertices) to make the sum in (2.1) small enough. This is because a decay of order $(1/\sqrt{d})^{|V(\alpha)|}$ fails to counterbalance the number of possible $\alpha\subseteq K_{n,m}$ already if $\alpha$ has only $n^{\varepsilon}$ vertices.
One possible approach for dealing with this issue (as also suggested in [3, Remark 1]) is to condition on a subset of the latent information given by one side of the bipartition. While this is indeed an important aspect of our approach, the arguments given in [3] would still not yield an exponent that is strong enough for our purposes. Moreover, as we will see later, we do not only need a bound on $\mathbb{E}\left[\textsc{Sw}(\alpha)\right]$, but we also need to explicitly control how exactly this quantity depends on the latent vectors we condition on, which seems to be out of reach of the techniques considered so far.